soybean meal is 16% protein, cornmeal is 8% protein. how many pounds of each should be mixed together in order to get 320-lb mixture that is 14% protein?
How many pounds of cornmeal?
How many pounds of soybean?

Answer by checkley77(12844)

.16x+.08(320-x)=.14*320
.16x+25.60-.08x=44.8
.08x=44.8-25.60
.08x=19.2
x=19.2/.08
x=240 lbs. of soybean are used.
320-240=80 lbs. of cornmeal is used.
Proof:
.16*240+.08*80=.14*320
38.4+6.4=44.8
44.8=44.8